OpenStudy (anonymous):

g(x)=frac(x-3){x(x+1)}

OpenStudy (anonymous):

\[g(x)=\frac{x-3}{x(x+1)}\]

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

yes find the vertical asymptotes if any

OpenStudy (anonymous):

set denominator = 0 and solve:\[x(x+1)=0\] \[x=0\] or \[x=-1\] so there are two.

OpenStudy (anonymous):

need the horizontal one as well? degree of numerator is 1,degree of denominator is 2, and since 2 > 1 it is \[y=0\]

OpenStudy (amistre64):

gotta make sure that the bottom doesnt cancel the top; if it does, that would be a hole; but this problem is good....

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\log_{c}t+\log_{c}s \]

OpenStudy (amistre64):

logc(ts)

OpenStudy (anonymous):

\[log_c(ts)\]

OpenStudy (anonymous):

hello again.

OpenStudy (anonymous):

oh so the bases never changes when re-writing??

OpenStudy (amistre64):

not in that case :)

OpenStudy (anonymous):

if they are different you cannot do it.

OpenStudy (anonymous):

k thanks ...was always confused about that..lol teachers dont kno what they are talking about..lol

OpenStudy (anonymous):

typical

OpenStudy (anonymous):

pre-calc class?

OpenStudy (anonymous):

sort of its a new thing in orlando ..its a college readiness class that is basically preparing you to take pre-calc in college

OpenStudy (anonymous):

\[\frac{2x}5=\frac{x}3+2;solve\]

OpenStudy (toxicsugar22):

i amristre can u help me after her

OpenStudy (amistre64):

orlando eh... im over near brooksville

OpenStudy (amistre64):

perhaps :) hypers are easy

OpenStudy (toxicsugar22):

amristre can u help me after her

OpenStudy (amistre64):

*15 all of it.... 6x = 5x +30 x = 30 maybe?

OpenStudy (amistre64):

dbl check

OpenStudy (toxicsugar22):

amristre can u help me after

OpenStudy (amistre64):

60/5 = 30/3 +2 6 = 10+2 .... gonna have to see what I did lol

OpenStudy (amistre64):

i can yes...

OpenStudy (amistre64):

2x x --- = --- + 2 5 3 2x = 5x/3 +10 2x = 5x+30/3 6x = 5x +30 x = 30 ..... 2(30) 30 ---- = --- + 2 5 3 2(6) = 10 + 2 12 = 12 ; well it worked that time lol.. my dbl check must have been off

OpenStudy (amistre64):

60/5 = 12 lol.....

OpenStudy (anonymous):

answer is either -+30 or -+60..but thanks lol

OpenStudy (amistre64):

x = 30 fer sure

OpenStudy (anonymous):

ok her is my impossible problem...lol

OpenStudy (anonymous):

x^2+7x+10 times x^2+6x ---------- ----------- multiply x^2+8x+12 x^2+13x+40

OpenStudy (amistre64):

factor to see what cancels... 2,5,x,6 ------ 6,2,8,5 ,,x, --- = x/(x+8) if I see it right ,,8,

OpenStudy (anonymous):

brilliant!

OpenStudy (anonymous):

but can you walk me through it so i can no how to do it myself

OpenStudy (anonymous):

\[\log_{3}...x=5 \]

OpenStudy (anonymous):

log_{3} x=5